OpenStudy (anonymous):

Pleeaaassseee help ! :( I'm having trouble with solving systems of two equations in two variables. It makes me feel dumb :( can someone please explain this example question for me? Solve the system of equations using elimination. a. -2x-3y=5 -5x+3y=40

OpenStudy (phi):

First, do you know how to simplify -3y + 3y ?

OpenStudy (anonymous):

yes

OpenStudy (phi):

you should get 0

OpenStudy (anonymous):

okay, so i'd eliminate that first. then should i try to get it by itself in one equation and get x?

OpenStudy (phi):

The other idea you need is if you have an equation, and you add the *same value* to both sides it will stay an equation (in other words, the left side = right side)

OpenStudy (phi):

Here is a simple example 1+2 = 3 (this is an equation) add 4 to both sides 1+2+4 = 3+4 (this is still true. if we simplify we get 7=7, which is true)

OpenStudy (phi):

now for your problem. we are given -2x-3y=5 left side equals the right side we are also told -5x+3y=40 (another equation) Let's start with the first equation and add 40 to both sides. (It will stay an equation right?) -2x-3y + 40 =5+40 now use the second equation, which says that 40 is the something as -5x+3y (that is what -5x+3y=40 means, right?) -2x-3y + -5x+3y = 5+40 simplify we know -3y +3y is 0 -2x -5x + 0 = 45 (and 5+40 is 45) now add -2x and -5x what do you get ?

OpenStudy (anonymous):

-7x

OpenStudy (phi):

yes, and we should always write out the full equation -7x= 45 divide both sides by -7

OpenStudy (anonymous):

x=6.4 ish? or 45/7

OpenStudy (phi):

you write \[ \frac{-7 \cdot x}{-7} = \frac{45}{-7} \] the -7/-7 on the left side is 1 \[ x = \frac{45}{-7} \] or - 45/7 so you are correct, except it's negative.

OpenStudy (anonymous):

opps ! yes sorry i forgot to add the negative sign ! Thanks for correcting me & thanks for helping !

OpenStudy (anonymous):

what it one thing i should always try to do first when trying to solve the equation though?

OpenStudy (phi):

I made it look complicated. Here's how people do it: Add the 2 equations: -2x-3y=5 -5x+3y=40 ----------- -7x = 45 now solve for x. That is the "short-way" to do what I did up above.

OpenStudy (phi):

Now you might be thinking... it was *awfully convenient* that the -3y + 3y=0 because otherwise, we would still have a "y" to confuse things.

OpenStudy (anonymous):

oh okay! my teacher didn't really explain this very well, that looks simple. If there is still a "y" , what would u do with it?

OpenStudy (phi):

if there is still a y, we are in trouble.

OpenStudy (anonymous):

Lol. So in this type of equation, there should only be one variable?

OpenStudy (anonymous):

*one variable left

OpenStudy (phi):

which is why we *make it so the y's go away* example: -2x -1y=5 -5x+3y=40 now if do our trick, we still have x's and y's and that is not good. but here is the idea (brilliant!) multiply the top equation (both sides) by 3 3 times the left side and 3 times the right side (it's still an equation) 3(-2x -1y) = 3*5 distribute the 3: -6x -3y = 15 now we have -6x -3y = 15 -5x+3y=40 and when we add the two equations, the y's go away.

OpenStudy (anonymous):

oh okay! understandable. In the equations before this, using substitution, there was 2 answers; one for x and one for y . Now, using elimination there should only be one, correct?

OpenStudy (phi):

there are two unknowns (the x and the y), and a complete solution has a number for both. but once you know what x is, we use either of the original equations, replace x with its number, and solve for y.

OpenStudy (anonymous):

okay! Thank You!

OpenStudy (phi):

If you want to see some examples, watch Khan's videos. For example, here's a 3 minute example: http://www.khanacademy.org/math/algebra/systems-of-eq-and-ineq/solving-systems-addition-elimination/v/addition-elimination-method-1